# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v2

@@
Contextual Stochastic Optimization for Omnichannel Multi-Courier Order Fulfillment Under Delivery Time Uncertainty
@@
NSF AI Institute for Advances in Optimization (AI4OPT), Georgia Institute of Technology, Atlanta, GA
@@ 1.1 Contributions
To address the challenge of observational data, this paper introduces a generalized CSO framework addressing these real-world complexities. The novel CSO framework is applied to a sophisticated stochastic omnichannel order fulfillment optimization problem with multiple carrier options and delivery time uncertainties. In collaborating with a major US online retailer, the framework is rigorously evaluated on an industrial order fulfillment dataset featuring a vast, intricate fulfillment network. This application demonstrates the framework capabilities and its potential impact on large-scale, real-world optimization challenges. The contributions of the paper can be summarized as follows.
@@
    Generic CSO framework for observational data: The paper presents a novel, distribution‑agnostic contextual stochastic optimization (CSO) framework that converts problems with missing‑counterfactual observational data into tractable problems. A contextual distribution oracle is learned from partial observations, and the resulting problem is solved via two scalable solution paradigms—Contextual Sample Average Approximation (C‑SAA) and Contextual Robust Optimization (C‑RO).
@@ 3.1.1 Overview of the Current Practice
Currently, the industrial partner uses a sourcing engine that employs a straightforward greedy algorithm. For each unit in an order, the sourcing engine sorts the eligible location-carrier pairs by the estimated shipping costs and transit times.
@@ 3.2 The Nominal Optimization Problem
This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem. A comprehensive list of the notations presented in this section is provided in Appendix for reference.
@@ 3.2.2 Objective Function
The objective of the optimization problem for an order is to minimize the overall fulfillment costs while ensuring timely delivery. The fulfillment costs for an order include the fixed sourcing (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair. Delivery timeliness is measured by deviations from the desired delivery date, whether early or late. To encourage timeliness, the objective imposes penalties for both early and late deliveries.
The objective function is formally defined as follows. Let $c_{ik\ell}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$. For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed sourcing cost. For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center. Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
@@
$g(\mathbf{z},\mathbf{d})=\begin{cases}\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}[c_{\ell}^{fixed}+(1-\beta_{k})c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-}]z_{ik\ell}&\text{if }\sum_{i\in\mathcal{I}}z_{ik\ell}\geq 2,\\ \sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-})z_{ik\ell}&\text{otherwise},\end{cases}$ (1)
@@ 3.3 The Contextual Stochastic Order Fulfillment Problem
(CSOFP) $\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (6)
@@
4 Methodology: A Generic CSO Framework for Observational Data
@@
This section lays out a generic contextual stochastic optimization (CSO) framework that does exactly that. In Section 4.1, the abstract CSO problem (denoted by ${\cal P}$) is defined. Section 4.2 shows how to construct a contextual distribution oracle $\mathcal{M}(\mathbf{s})$ from partial observations. Sections 4.3 and 4.4 introduce two solution paradigms—Contextual Sample Average Approximation and Contextual Robust Optimization—which use $\mathcal{M}(\mathbf{s})$ as input. Finally, Section 4.6 maps all of the above back to the omnichannel multi-courier CSOFP introduced in Section 3.
@@ 4.1 Generic Contextual Stochastic Optimization (CSO) Problem
Problem ${\cal P}$ consists of fulfilling a set of incoming requests sequentially, each of which specifies demands for a subset of products. Let $\mathcal{I}$ be the set of products, $\mathcal{V}$ the set of selection options (e.g., sourcing or transport modes), and $\mathcal{Z}\subseteq\mathbb{Z}_{+}^{|\mathcal{I}|\times|\mathcal{V}|}$ the feasible assignment set. Each request arrives with known demand quantities $\{q_{i}:i\in\mathcal{I}\}$ at decision time. The decision variable $\mathbf{z}=(z_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\in\mathcal{Z}$ assigns quantity $z_{iv}$ of product $i$ to option $v$.
Let $\tilde{\mathbf{c}}=(\tilde{c}_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\sim\mathbb{P}$ denote the random cost vector. Assume that $\tilde{\mathbf{c}}$ is element-wise positive and the cost function $g(\mathbf{z},\tilde{\mathbf{c}})$ is piecewise linear in $\mathbf{z}$. Introduce, for $j=0,1,\dots,n$, the breakpoint vectors
@@
$[\mathbf{b}^{j-1},\,\mathbf{b}^{j}]=\bigl{\{}\mathbf{z}:b^{\,j-1}_{iv}\leq z_{iv}\leq b^{j}_{iv}\;\forall\,i\in\mathcal{I},\;v\in\mathcal{V}\bigr{\}}.$
@@
$\min_{\mathbf{z}\in\mathcal{Z}}\;\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}}\bigl{[}g(\mathbf{z},\tilde{\mathbf{c}})\bigr{]}.$ (8)
@@
$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\mathbb{P}(\mathbf{c}\mid\mathbf{s})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (9)
@@
$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{c}}\sim\Pi_{v\in\mathcal{V}}\mathbb{P}_{v}(\mathbf{c}_{v}\mid\mathbf{s}_{v})}[g(\mathbf{z},\mathbf{\tilde{c}})].$ (10)
@@ 4.4.2 Contextual Budgeted Interval Uncertainty Set
[…]
@@ 4.5 Performance Guarantees on C-SAA & C-RO
[…]
@@ 4.6 Application to the Omnichannel Multi-Courier CSOFP
The generic CSO framework applies directly to CSOFP via the following identifications:
@@
$\min_{\mathbf{z}\in\mathcal{Z}}\mathbb{E}_{\tilde{\mathbf{d}}\sim\Pi_{k,\ell}\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})}[g(\mathbf{z},\tilde{\mathbf{d}})].$ (17)
To learn each marginal distribution $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$, partition the historical dataset into
$\mathbb{D}_{k\ell}=\bigl{\{}(\mathbf{s}_{k\ell,m},\,d_{k\ell,m})\bigr{\}}_{m=1}^{M_{k\ell}},$
and train a per‑carrier model $\mathcal{M}_{k}$ on $\mathbb{D}_{k}=\bigcup_{\ell\in\mathcal{L}}\mathbb{D}_{k\ell}$. The full contextual distribution oracle is $\mathcal{M}(\mathbf{s})=\bigl{(}\mathcal{M}_{k}(\mathbf{s}_{k\ell})\bigr{)}_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$
This per‑carrier decomposition simplifies learning, generalizes to new locations, and outperforms a single global model according to preliminary experiments.
@@ 4.6.1 Applying C-SAA and C-RO
Computing the expectation in $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathcal{M}(\mathbf{s})}\bigl{[}g(\mathbf{z},\tilde{\mathbf{d}})\bigr{]}$ directly would require either enumerating all $C^{|\mathcal{K}|\times|\mathcal{L}|}$ deviation vectors or numerically integrating a continuous CDF—both computationally intractable. Instead, the two solution paradigms use the contextual distribution oracle $\mathcal{M}(\mathbf{s})$ as follows:
@@
Both approaches avoid the exponential support growth while remaining agnostic to the internal structure of $\mathcal{M}$. A full C‑RO‑B reformulation is provided in the Appendix.
@@ 5.2 Probabilistic Multi-Class Classification
Multinomial logistic regression (MLR) and classification tree-based models are considered to solve MC-CLF. Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, the negative of the log loss function. The log loss is defined as $-\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
@@
To enhance predictive power, ensemble learning methods such as bagging and gradient boosting are employed, through the use of random forests (RF) and gradient boosted trees (GBT). The details for their calculation of the class probabilities are available in Appendix.
@@ 5.2.1 Probability Calibration
In this paper, a $k$-fold cross‐validation procedure is employed: for each fold $i$, the classifier is trained on its training split, raw scores are obtained on the held‐out calibration split, and an isotonic regression model $g^{(i)}$ is fit to those scores. On unseen test instance with covariates $\mathbf{s}$, each fold’s classifier and corresponding $g^{(i)}$ yield calibrated probabilities $\tilde{p}_{c}^{(i)}(\mathbf{s})$, which are then averaged across all $k$ folds to produce the final estimate $\tilde{p}_{c}(\mathbf{s})=\frac{1}{k}\sum_{i=1}^{k}\tilde{p}_{c}^{(i)}(\mathbf{s})\,.$
@@ 5.3 Tree-Based Quantile Regression
where $\hat{d}$ denotes the model’s predicted $q$‑th quantile. Collecting these quantile estimates across $q$ yields a stepwise approximation of the CDF. However, estimating each quantile independently can violate the ordering $\hat{d}_{(q_{1})}\leq\hat{d}_{(q_{2})}$ for $q_{1}<q_{2}$, resulting in a non‑monotonic—and thus invalid—CDF.
@@
[…]
@@ 6.3 Baseline Models
    Pure Cost-Driven Greedy Heuristic (Greedy): Mimics the industrial partner’s sourcing engine by selecting, for each unit of an SKU, the carrier–location pair that minimizes the fulfillment cost. Formally, for each order it solves $\min_{\mathbf{z}\in\mathcal{Z}}\{\sum_{i,k,\ell}(c_{\ell}^{fixed}+c_{ik\ell}^{ship})z_{ik\ell}\}.$
@@
    Point Predict–Then–Optimize (PTO): Trains per-carrier regression models $h_{k}(\cdot)$ to estimate the conditional mean deviations for each $(k,\ell)\in\mathcal{K}\times\mathcal{L}$, producing the point‐prediction vector $\hat{\mathbf{d}}=\{h_{k}(\mathbf{s}_{k\ell})\}_{k\in\mathcal{K},\ell\in\mathcal{L}}$. It then solves the deterministic problem $\min_{\bm{z}\in\mathcal{Z}}g\bigl{(}\bm{z},\hat{\mathbf{d}}\bigr{)}.$
    This method is effective only if the objective function $g(\mathbf{z},\tilde{\mathbf{d}})$ is linear in $\tilde{\mathbf{d}}$, in which case estimating the conditional distribution simplifies to estimating the conditional mean. This is because, by linearity of expectation, $\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[g(\mathbf{z}_{o},\tilde{\mathbf{d}})]=g(\mathbf{z},\mathbb{E}_{\tilde{\mathbf{d}}\sim\mathbb{P}(\mathbf{d}\mid\mathbf{s})}[\tilde{\mathbf{d}}])$. However, $g(\bm{z},\tilde{\mathbf{d}})$ is nonlinear due to asymmetric penalty costs of the deviations, hence replacing the distribution by its mean may bias the solution. For each carrier, the regression model with the lowest cross‐validated mean squared error is selected (see Appendix for details).
@@ 6.4.2 Simulation Framework
One of the difficulties in this application is that the ground truth is not available for options that were not selected. To overcome this limitation, a simulation environment was built to generate synthetic “ground truth” when an algorithm selects an option. Specifically, delivery deviations were drawn from a multinomial distribution, with probabilities estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders from the last week of the dataset (testing period). Obviously, this data was not accessible to the algorithm and only used by the simulator to replicate real-world conditions. The training and calibration procedures of the simulator closely followed the methodology outlined in Section 5.2. An evaluation of the simulator is provided in Appendix.

### v3

# MSOM-2024-1328.R2

@@
Contextual Stochastic Optimization for
Omnichannel Multi-Courier Order Fulfillment Under Delivery Time Uncertainty
@@
[…]
@@ 1.1 Contributions
To address the challenge of observational data, this paper introduces a generalized CSO framework addressing these real-world complexities. The novel CSO framework is applied to a sophisticated stochastic omnichannel order fulfillment optimization problem with multiple carrier options and delivery time uncertainties. In collaborating with a major US online retailer, the framework is rigorously evaluated on an industrial order fulfillment dataset featuring a vast, intricate fulfillment network. This application demonstrates the framework capabilities and its potential impact on large-scale, real-world optimization challenges. Based on the significant performance improvements demonstrated in this study, the industrial partner is actively considering a pilot program to integrate key elements of the proposed framework into their production fulfillment optimization engine. The contributions of the paper can be summarized as follows.
@@
    A modular and scalable CSO framework for observational data: The paper presents a novel, distribution‑agnostic contextual stochastic optimization (CSO) framework that converts problems with missing‑counterfactual observational data into tractable problems. A contextual distribution oracle is learned from partial observations, and the resulting problem is solved via two scalable solution paradigms—Contextual Sample Average Approximation (C‑SAA) and Contextual Robust Optimization (C‑RO). This features a tractable formulation for assignment-type problems with piecewise linear objectives.
@@ 3.1.1 Overview of the Current Practice
Currently, the industrial partner uses a fulfillment optimization engine that employs a straightforward greedy algorithm. For each unit in an order, the fulfillment optimization engine sorts the eligible location-carrier pairs by the estimated shipping costs and transit times.
@@ 3.2 The Nominal Optimization Problem
This section introduces the nominal MILP model developed for optimizing the omnichannel multi-courier order fulfillment problem. A comprehensive list of the notations presented in this section is provided in Appendix A for reference.
@@ 3.2.2 Objective Function
The objective of the optimization problem for an order is to minimize the overall fulfillment costs while ensuring timely delivery. The fulfillment costs for an order include the fixed fulfillment (non-parcel) costs and the shipping (parcel) costs, with consolidation discounts applied when multiple units are assigned to the same carrier-location pair. Delivery timeliness is measured by deviations from the desired delivery date, whether early or late. To encourage timeliness, the objective imposes penalties for both early and late deliveries.
The objective function is formally defined as follows. Let $c_{ik\ell}^{ship}>0$ be the per-unit shipping cost when using location $\ell$ and carrier $k$ to fulfill SKU $i$. For each location $\ell$, let $c_{\ell}^{fixed}>0$ be the per-unit fixed fulfillment cost. For each carrier $k$, let $\beta_{k}\in(0,1)$ be the discount percentage applied to shipping costs when multiple units are shipped using carrier $k$ from the same fulfillment center. Let $\gamma^{+}\geq 0$ and $\gamma^{-}\geq 0$ be constants that convert late and early delivery penalties into per-unit costs, respectively.
@@
[…]
@@ 3.3 The Contextual Stochastic Order Fulfillment Problem
[…]
@@
4 Methodology: A CSO Framework for Observational Data
@@
This section lays out a contextual stochastic optimization (CSO) framework that does exactly that. In Section 4.1, a generic CSO problem (denoted by $\cal P$}) is defined. Section 4.2 shows how to construct a contextual distribution oracle $\mathcal{M}(\mathbf{s})$ from partial observations. Sections 4.3 and 4.4 introduce two solution paradigms—Contextual Sample Average Approximation and Contextual Robust Optimization—which use $\mathcal{M}(\mathbf{s})$ as input. Finally, Section 4.6 maps all of the above back to the omnichannel multi-courier CSOFP introduced in Section 3. Figure 2 provides a schematic overview of the proposed framework, illustrating the flow from historical observational data through the learning and optimization stages to produce a final decision.
Figure 2: A Schematic Overview of the CSO Framework for Observational Data.
@@ 4.1 Generic Contextual Stochastic Optimization (CSO) Problem
Problem $\cal P$} consists of fulfilling a set of incoming requests sequentially, each of which specifies demands for a subset of products. Let $\mathcal{I}$ be the set of products, $\mathcal{V}$ the set of selection options (e.g., fulfillment or transport modes), and $\mathcal{Z}\subseteq\mathbb{Z}_{+}^{|\mathcal{I}|\times|\mathcal{V}|}$ the feasible assignment set. Each request arrives with known demand quantities $\{q_{i}:i\in\mathcal{I}\}$ at decision time. The decision variable $\mathbf{z}=(z_{iv})_{i\in\mathcal{I},\;v\in\mathcal{V}}\in\mathcal{Z}$ assigns quantity $z_{iv}$ of product $i$ to option $v$.
@@
$\mathbf{b}^{j}=\bigl(b^{j}_{iv}\bigr)_{i\in\mathcal{I},\;v\in\mathcal{V}},~\mathbf{b}^{0}<\mathbf{b}^{1}<\cdots<\mathbf{b}^{n}\quad(\text{component-wise}),$
@@
$\mathbf{b}^{j-1},\,\mathbf{b}^{j}]=\bigl\{\mathbf{z}:b^{\,j-1}_{iv}\leq z_{iv}\leq b^{j}_{iv}\;\forall\,i\in\mathcal{I},\;v\in\mathcal{V}\bigr\}.$
@@
[…]
@@
[…]
@@
[…]
@@ 4.4.2 Contextual Budgeted Interval Uncertainty Set
$\text{(C-RO-B)}\quad\min_{\mathbf{z}\in\mathcal{Z},y,w,\pi,\lambda}$ $(\sum_{i,v,j}w_{iv}^{j}+\pi B+\sum_{i,v}\lambda_{iv})$ (16)
@@ 4.5 Performance Guarantees on C-SAA & C-RO
$N\;\geq\;2|\mathcal{V}||\mathcal{I}|\;+\;\frac{2|\mathcal{V}||\mathcal{I}|}{\epsilon}\ln\!\Bigl(\frac{2}{\epsilon}\Bigr)\;+\;\frac{2}{\epsilon}\ln\!\Bigl(\frac{1}{\delta}\Bigr)$
@@ 4.6 Application to the Omnichannel Multi-Courier CSOFP
The CSO framework applies directly to CSOFP via the following identifications:
@@
[…]
To learn each marginal distribution $\mathbb{P}_{k\ell}(d_{k\ell}\mid\mathbf{s}_{k\ell})$, the historical dataset is first partitioned into $\mathbb{D}_{k\ell}=\bigl\{(\mathbf{s}_{k\ell,m},\,d_{k\ell,m})\bigr\}_{m=1}^{M_{k\ell}}.$ This is followed by training a per‑carrier model $\mathcal{M}_{k}$ on the dataset $\mathbb{D}_{k}=\bigcup_{\ell\in\mathcal{L}}\mathbb{D}_{k\ell}$. This per-carrier approach is a practical adaptation of the generic per-option framework from Section 4.2. Rather than training a distinct model for each individual location-carrier pair (which can be inefficient and suffer from data sparsity), the data is pooled to train a single, more robust model for each carrier. This decomposition simplifies learning, allows the model to generalize to new locations, and was found to outperform a single global model in preliminary experiments. The full contextual distribution oracle is then assembled as $\mathcal{M}(\mathbf{s})=\bigl(\mathcal{M}_{k}(\mathbf{s}_{k\ell})\bigr)_{k\in\mathcal{K},\,\ell\in\mathcal{L}}.$
@@ 4.6.1 Applying C-SAA and C-RO
[…]
@@
Both approaches avoid the exponential support growth while remaining agnostic to the internal structure of $\mathcal{M}$. A full C‑RO‑B reformulation is provided in the Appendix F.
@@ 5.2 Probabilistic Multi-Class Classification
Multinomial logistic regression (MLR) and classification tree-based models are considered to solve MC-CLF. Specifically, MLR is a classic classifier that aims to estimate the class probabilities by maximizing the log-likelihood function, the negative of the log loss function. The log loss is defined as $\sum_{m=1}^{M}\sum_{c=1}^{C}\mathbf{1}\{d_{m}=\xi_{c}\}\log(\hat{p}_{c}(\mathbf{s}_{m})).$ On the other hand, classification trees use a nonparametric approach that recursively partitions the data into subsets based on feature values in a top-down manner, resulting in a tree-like model of decisions.
@@
To enhance predictive power, ensemble learning methods such as bagging and gradient boosting are employed, through the use of random forests (RF) and gradient boosted trees (GBT). The details for their calculation of the class probabilities are available in Appendix G.
@@ 5.2.1 Probability Calibration
In this paper, a $k$-fold cross‐validation procedure is employed: for each fold $i$, the classifier is trained on its training split, raw scores are obtained on the held‐out calibration split, and an isotonic regression model $g^{(i)}$ is fit to those scores. On unseen test instance with covariates $\mathbf{s}$, each fold’s classifier and corresponding $g^{(i)}$ yield calibrated probabilities $\tilde{p}_{c}^{(i)}(\mathbf{s})$, which are then averaged across all $k$ folds to produce the final estimate $\tilde{p}_{c}(\mathbf{s})=\frac{1}{k}\sum_{i=1}^{k}\tilde{p}_{c}^{(i)}(\mathbf{s})\,.$ Since per-class calibration does not guarantee the final probabilities will sum to one, a normalization step is applied. The vector of calibrated class probabilities $\bigl(\tilde{p}_{c}(\mathbf{s})\bigr)_{c=1}^{C}$ is divided by its sum to yield a valid multi-class probability distribution.
@@ 5.3 Tree-Based Quantile Regression
where $\hat{d}$ denotes the model’s predicted qq‑th quantile. Collecting these quantile estimates across $q$ yields a stepwise approximation of the CDF. However, estimating each quantile independently can violate the ordering $\hat{d}_{(q_{1})}\leq\hat{d}_{(q_{2})}$ for $q_{1}<q_{2}$, resulting in a non‑monotonic—and thus invalid—CDF.
@@
$\text{CRPS}(\hat{F},d,\mathbf{s})=\int_{-\infty}^{\infty}(\hat{F}(u\mid\mathbf{s})-\mathbf{1}\{u\geq d\})^{2}du.$ (22)
@@ 6.3 Baseline Models
    Pure Cost-Driven Greedy Heuristic (Greedy): Mimics the industrial partner’s fulfillment optimization engine by selecting, for each unit of an SKU, the carrier–location pair that minimizes the fulfillment cost. Formally, for each order it solves $\min_{\mathbf{z}\in\mathcal{Z}}\{\sum_{i,k,\ell}(c_{\ell}^{fixed}+c_{ik\ell}^{ship})z_{ik\ell}\}.$
@@
[…]
@@ 6.4.2 Simulation Framework
One of the difficulties in this application is that the ground truth is not available for options that were not selected. To overcome this limitation, a simulation environment was built to generate synthetic “ground truth” when an algorithm selects an option. Specifically, delivery deviations were drawn from a multinomial distribution, with probabilities estimated by fitting a fine-tuned, calibrated multi-class classification model based on orders from the last week of the dataset (testing period). Obviously, this data was not accessible to the algorithm and only used by the simulator to replicate real-world conditions. The training and calibration procedures of the simulator closely followed the methodology outlined in Section 5.2. An evaluation of the simulator is provided in Appendix J.